# Can GitHub Issues Help in App Review Classifications?

### v2

@@ 4.3 Evaluation Measures
$precision=TP/(TP+FP)$ (1)
$recall=TP/(TP+FN)$ (2)
$F1-score=2*precision*recall/(precision+recall)$ (3)
@@
As our selected primary datasets exhibit an imbalanced distribution, we employed a 5-fold cross-validation approach to mitigate potential result bias. Consequently, the results obtained represent the average metric values across each fold. To ensure consistent distribution within each fold, we used the StratifiedKFold function [59] from the sklearn library for fold selection [22].
5 Experiment Design and Implications
5.1 RQ1: Influences of Augmentation Methods on the Model Performance
@@ 5.1.1 Results
Table 12: Results of Bug Report Classification on Truth Dataset Using Primary (Rev) and Augmented (Rev+(S)I, Rev+(Sim)I, Rev+(S+Sim)I, and Rev+(Rand)I) Datasets in RQ1 (Rev=Reviews, I=Issues, S=Same, Sim=Similar, Rand=Random).
@@
Table 13: Results of Feature Request Classification on Truth Dataset Using Primary (Rev) and Augmented (Rev+(S)I, Rev+(Sim)I, Rev+(S+Sim)I, and Rev+(Rand)I) Datasets in RQ1 (Rev=Reviews, I=Issues, S=Same, Sim=Similar, Rand=Random).
@@
Since reviews represent the user requirements, analysis of them can affect software maintenance decisions to promote the user’s satisfaction and the app’s success [60]. Recall can be more important than precision from the developer’s point of view in the review classification model related to certain situations, especially for identifying bug reports, as developers probably prefer not to miss them [1]. This is because a high recall ensures that the model identifies most, if not all, actual bug reports. While a high precision ensures that the model mainly identifies true bugs and avoids irrelevant reviews, missing even a single bug report can be more detrimental for developers. Unfixed bugs can lead to crashes, errors, and user frustration, whereas developers can quickly dismiss false positives. Generally, obtained results show that although the average precision decreases in some augmented models related to bug reports, the F1-score increases due to the significant recall improvement. Therefore, models trained on augmented datasets identify a higher percentage of actual bug report reviews that are helpful for developers. For example, in the bug report, the recall of the baseline model is between 50.4 to 67.2 and, on average, 61.4. In our models, recall is between 47.9 to 83.3; on average, it is between 61.4 to 70.6. Additionally, the data augmentation makes the training dataset more diverse and provides more samples for different types of reviews. Hence, trained models on augmented datasets are less biased to specific types, which increases recall.
@@
On the other hand, in the bug report, the average F1-score associated with Rev+(Sim)I, 69.5, is better than Rev+(S)I, 66.1. However, in the feature request, we observe the opposite results, where Rev+(S)I outperforms Rev+(Sim)I with an average F1-score of 66.2 compared to 64.3. The reason for this difference could be that bug reports for similar apps exhibit a higher degree of similarity to each other [61]. Encountered bugs usually include common problems such as crashing, making errors, freezing screens, etc., which can be found in similar apps. Meanwhile, due to the wide range of available apps in GooglePlay, app features are highly varied, which makes it challenging to identify requests for them using our limited dataset. Each application has specific features that are mentioned in issues or reviews. So, bug reports in the similar apps can have more similarity to each other than feature requests. Therefore, the better performance of the Rev+(Sim)I compared to the Rev+(S)I for bug reports and the opposite results in the feature requests can be rationalized.
@@
5.2 RQ2: Influences of Dataset Augmentation on Increasing the Generalizability of Model
In RQ2, we investigate the impact of augmenting primary datasets on increasing the generalizability of the trained classifiers. Devine et al. [7] demonstrated that models trained on seven labeled datasets do not perform well in predicting unseen ones. Therefore, this experiment demonstrates that augmenting review datasets using auxiliary datasets extracted from labeled issues can contribute to the enhancement of performance in trained classifiers. To this end, we use five primary datasets (Table 10) for training baseline and augmented models. Devine et al. [7] introduced two types of models, namely “Single-Dataset” and “Leave-One-Out (LOO)”, for comparison purposes. Their study revealed that the LOO models yield superior results. In this experiment, we also train these two model types using the primary and augmented datasets and compare their results:
  1. 1.
@@ 5.2.1 Results
Table 14: Results of Single-Dataset and LOO Models for Bug Report Classification Using Primary and Augmented Datasets in RQ2 (PD=Primary Dataset, AD=Auxiliary Dataset).
@@
Table 15: Results of Single-Dataset and LOO Models for Feature Request Classification Using Primary and Augmented Datasets in RQ2 (PD=Primary Dataset, AD=Auxiliary Dataset).
@@
Devine et al. [7] investigate the performance of review classifiers in predicting new datasets that have not seen any review from them in the training dataset. They demonstrate that models trained on LOO (combined multiple datasets) outperformed Single-Dataset models. In this experiment, LOO models do not necessarily provide the best results but are close to maximum scores. In the bug report, models with the highest F1-scores in order related to augmented PD3 (67.8), augmented PD4 (67.6), baseline PD3 (67.3), and baseline LOO (67.1). In the feature request, models with the highest F1-score in order related to baseline PD4 (50.9), baseline LOO (49.1), and jointly augmented LOO and PD3 (48.8). So, models that augment with the Between-App method can outperform LOO models.
@@
5.3 RQ3: Influences of Volume of Auxiliary Dataset on Model Performance
@@ 5.3.1 Results
Table 16: Results of Bug Report Classification Using Augmented Datasets with Varied Volumes in RQ3 (AD=Auxiliary Dataset).
@@
Table 17: Results of Feature Request Classification Using Augmented Datasets with Varied Volumes in RQ3 (AD=Auxiliary Dataset).
@@
Figure 9: The Average F1-score and Recall for the Bug Report/Feature Request Classifiers Trained on Datasets Augmented with Various Volume Ratios.
@@
5.4 RQ4: Influences of Dataset Augmentation on Transformer-Based Models
In this research question, our goal is to explore the impact of dataset augmentation on the performance of classifiers trained using various transformer-based algorithms. We discuss the results of review classifiers based on four models BERT [18], RoBERTa [19], ALBERT [21], and DistilBERT [20]. RoBERTa [62] stands as “Robustly-Optimized BERT” that performs better than BERT [63] by applying modifications in training and masking strategies. Also, ALBERT [64] and DistilBERT [65] use techniques to provide faster and lighter models without significant performance degradation [66]. Since transformer-based models outperformed traditional algorithms for our task in previous studies [22, 46], we do not consider them in this experiment. The results of this experiment will facilitate the selection of the appropriate algorithm for training binary review classifiers.
5.4.1 Results
@@
Table 18: Results of Review Classification Using Primary (LOO) and Augmented Datasets with Different Transformer-Based Models in RQ4 (*minimum time)
@@
5.5 Discussion
In this paper, our goal is to extract valuable information from the labeled issues and use them to enrich the training dataset. Our experiments show that using the processed issues as an auxiliary dataset result in an improved review classifier, notably in terms of F1-score and recall metrics. User reviews, especially those related to the bug report, can help developers in maintenance decisions [31]. Therefore, losing a smaller percentage of user reviews related to app maintenance enhances the reliability of the review classifier. Also, since reviews may change over time [8], using processed issues as part of the training dataset makes it easier and faster to update it.
@@ 5.5 Discussion
5.5.1 Training App-Specific Classifiers
@@
5.5.2 Training General-Purpose Classifiers
@@ 6 Related Work
6.1 Approaches Based on App Store Reviews
@@
Some previous research analyzes user reviews based on Natural Language Processing (NLP) and Information Retrieval (IR) techniques. As one of the first research in user review analysis, Iacob et al. [67] propose the MARA tool. They conduct manual analysis and define 237 linguistic rules for extracting feature requests. They rank identified requests by their length and frequency. Guzman et al. [68] automatically analyze user reviews and identify app features at two levels. They filter out the extracted collocations to discover low-level features and compute sentiment scores for them. These features are then organized into groups using topic modeling techniques. Gao et al. [69] propose an approach that automatically extracts and prioritizes the main topics derived from user reviews. In addition to the information provided by users, they also use the relevance of the topics to rank them. Also, they visualize and analyze topic dynamics over time. While these unsupervised approaches aid in extracting valuable insights from user reviews, it’s important to note that supervised algorithms can provide better performance in identifying reviews related to predefined classes.
@@
Many studies used traditional learning algorithms to train user review classifiers. Chen et al. [5] present AR-Miner, an approach to identify informative user reviews using the EMNB algorithm and group them by applying the topic modeling. They provide a scheme for ranking generated groups and their instances. Villarroel et al. [4] [34] present the CLAP tool that helps developers plan the next release. They classify user reviews in the bug report, suggestion for new feature, and other classes using their constituent words, N-grams, and ratings. Also, they cluster the reviews related to the two first classes, which are of developer interest. Then, clusters were prioritized to recommend the most important ones by defining several independent features. Maalej et al. [41][1] classify reviews into four classes bug report, feature request, user experience, and rating. They employ this approach based on constituent words, collocations, general metadata, and sentiment scores derived from user reviews. Panichella et al. [30] classify review sentences using a combination of three features: constituent terms, grammatical patterns, and sentiment scores. They analyze developer discussions about maintenance tasks and adapt them with user review categories provided by Pagano et al. [29]. As a result, they select helpful categories from the developer’s perspective relevant to the app evolution. Guzman et al. [6] identify a taxonomy of seven categories (bug report, feature strength, feature shortcoming, user request, praise, complaint, and usage scenario) that provide helpful information for the app evolution. They represent the vectorized form of a user review by the tf-idf weighted vector and several numerical features (such as the word and character count). Also, they use four traditional algorithms and their ensembles to classify user reviews based on their taxonomy. Jha et al. [33] present an approach to classify and summarize helpful reviews. In this approach, reviews represent by semantic frames that are automatically extracted and more efficient than linguistic patterns obtained by manual analysis. Gu et al. [31] aim to summarize user reviews. First, they classify reviews into five categories, aspect evaluation, bug report, feature request, praise, and other, and use two types of textual features, lexical and structural. Also, they summarize the sentences of the aspect evaluation category by identifying and clustering the aspects and their related opinions and sentiments. This approach employs defined patterns to extract app aspects.
@@
Recent studies show that deep learning algorithms can outperform traditional algorithms in review classification. Aslam et al. [2] use a CNN-based network that considers the textual, non-textual and sentiment features extracted from app reviews as classifier inputs. Also, Stanik et al. [70] present an approach to classify user reviews and tweets in two languages, English and Italian. In addition to traditional models, they use a CNN-based network for classifier training. Due to the relatively small volume of labeled datasets, transfer learning can be helpful in review classification task [46]. Henao et al. [46] argue that BERT-based models outperform models based on traditional machine learning algorithms. Also, Araujo et al. [47] compare the various techniques available for review representation and demonstrate that the pre-trained transformer-based models achieve better results in review analysis tasks. Hadi et al. [22] conduct a more comprehensive study comparing pre-trained models and previous approaches for the review classification task in different settings. They focused their research on six labeled datasets from prior works and integrated them and four popular approaches. In addition, they pre-train new models specific to this domain using collected user reviews and show that using them results in better performance. So, we also use pre-trained models in our approach. In most previous works, predictive models have been trained and tested on one dataset. Devine et al. [7] discuss how a review classifier trained on one dataset performs well in predicting the intention of reviews related to other apps or datasets. This study uses seven datasets labeled in previous studies and shows that the results of model evaluations on unseen datasets cannot perform as well as the model evaluations on the same dataset. Also, they argue that merging labeled datasets can help overcome this generalizability problem. Our paper provides another solution by augmenting the labeled review dataset with processed issues that can result in better performance than the merged dataset.
6.2 Approaches Based on Other Sources of data
@@
Different sources, such as issue tracker systems, developer emails, and social platforms, contain valuable information for development and maintenance tasks [7]. Twitter is one of the sources that has been the focus of researchers. Guzman et al. [71] manually analyze tweets related to software applications (including desktop, mobile, or both) and identify 26 categories for their content. Also, they train traditional classifiers to predict the relevance of these categories with the various types of stakeholders. Nayebi et al. [13] classify tweets and user reviews into three categories: bug reports, feature requests, and others. Then, they extract the topics of each category and analyze their similarity in reviews and tweets. They found that analyzing tweets can provide developers with worthy information besides information extracted from reviews. Sorbo et al. [72] classify the information provided in the developer emails into six categories with manual analysis. They define linguistic patterns to identify the corresponding category. Shi et al. [73] present an approach to identify feature requests in developer emails. They categorize sentences using 81 fuzzy rules and find feature requests using sequential pattern mining. In these approaches, training and testing datasets are selected from the same source.
@@
Some studies use available data from one source to help analyze and extract information from another source. Huang et al. [74] examine whether the taxonomy introduced by Sorbo et al. [72], provided by analyzing the email contents, can be useful for categorizing GitHub Comments. They customize it and make it suitable for the analysis of GitHub comments. Also, they use a CNN-based neural network to categorize them automatically. Their results show that using classifiers trained on a dataset related to one source (developer email/issues) to predict a dataset related to another source (issues/developer email) does not provide better results than the classifiers that use data from the same source for training and prediction. Therefore, since data from different sources usually have a different nature, we consider the size of the auxiliary datasets to be smaller than the primary datasets in our approach to avoid the excessive biasing model to them. Stanik [70], Henao [46], and Devine [7] also use Twitter datasets in their studies. However, only the last one examines merging them with review datasets to improve the performance of trained models.
@@
Al-Safoury et al. [16] aggregate the information provided in user reviews and GitHub issues. They extract bug reports and feature requests from them and merge similar ones to create a complete change feature list by computing textual similarity. In our proposed approach, we process issues to adapt with app store reviews. Then, we use them to augment the labeled review dataset. As far as we know, GitHub issues have not been used in previous approaches to improve the performance of the review classification. Zhang et al. [17] utilized the similarity between user reviews and GitHub issues to categorize unlabeled issues. They applied the SURF tool [28] to identify app reviews related to problem discovery and feature requests and incorporated their similarity with unlabeled issues in the final similarity measure. In this study, our objective differs from their research, as we utilize labeled issues to assist in categorizing user reviews.
@@
Additionally, some studies augment the text of reviews to improve the accuracy of review prediction. Wang et al. [75] emphasize that using other sources can help prepare accurate classifiers. They extract feature words related to each category from app changelogs and use them to augment reviews. Liu et al. [76] also utilize app descriptions for this purpose. Our proposed approach differs from this research because we aim to augment the labeled dataset and not add additional information to the text of reviews.
7 Threats to Validity

### v3

@@ 4.3. Evaluation Measures
(1) $precision=TP/(TP+FP)$
(2) $recall=TP/(TP+FN)$
(3) $F1-score=2*precision*recall/(precision+recall)$
@@
As our selected primary datasets exhibit an imbalanced distribution, we employed a 5-fold cross-validation approach to mitigate potential result bias. Consequently, the results obtained represent the average metric values across each fold. To ensure consistent distribution within each fold, we used the StratifiedKFold function (Str, 2023) from the sklearn library for fold selection (Hadi and Fard, 2021).
5. Experiment Design and Implications
5.1. RQ1: Influences of Augmentation Methods on the Model Performance
@@ 5.1.1. Results
Table 12. Results of Bug Report Classification on Truth Dataset Using Primary (Rev) and Augmented (Rev+(S)I, Rev+(Sim)I, Rev+(S+Sim)I, and Rev+(Rand)I) Datasets in RQ1 (Rev=Reviews, I=Issues, S=Same, Sim=Similar, Rand=Random).
@@
Table 13. Results of Feature Request Classification on Truth Dataset Using Primary (Rev) and Augmented (Rev+(S)I, Rev+(Sim)I, Rev+(S+Sim)I, and Rev+(Rand)I) Datasets in RQ1 (Rev=Reviews, I=Issues, S=Same, Sim=Similar, Rand=Random).
@@
Since reviews represent the user requirements, analysis of them can affect software maintenance decisions to promote the user’s satisfaction and the app’s success (Palomba et al., 2015). Recall can be more important than precision from the developer’s point of view in the review classification model related to certain situations, especially for identifying bug reports, as developers probably prefer not to miss them (Maalej et al., 2016). This is because a high recall ensures that the model identifies most, if not all, actual bug reports. While a high precision ensures that the model mainly identifies true bugs and avoids irrelevant reviews, missing even a single bug report can be more detrimental for developers. Unfixed bugs can lead to crashes, errors, and user frustration, whereas developers can quickly dismiss false positives. Generally, obtained results show that although the average precision decreases in some augmented models related to bug reports, the F1-score increases due to the significant recall improvement. Therefore, models trained on augmented datasets identify a higher percentage of actual bug report reviews that are helpful for developers. For example, in the bug report, the recall of the baseline model is between 50.4 to 67.2 and, on average, 61.4. In our models, recall is between 47.9 to 83.3; on average, it is between 61.4 to 70.6. Additionally, the data augmentation makes the training dataset more diverse and provides more samples for different types of reviews. Hence, trained models on augmented datasets are less biased to specific types, which increases recall.
@@
On the other hand, in the bug report, the average F1-score associated with Rev+(Sim)I, 69.5, is better than Rev+(S)I, 66.1. However, in the feature request, we observe the opposite results, where Rev+(S)I outperforms Rev+(Sim)I with an average F1-score of 66.2 compared to 64.3. The reason for this difference could be that bug reports for similar apps exhibit a higher degree of similarity to each other (Tan and Li, 2020). Encountered bugs usually include common problems such as crashing, making errors, freezing screens, etc., which can be found in similar apps. Meanwhile, due to the wide range of available apps in Google Play, app features are highly varied, which makes it challenging to identify requests for them using our limited dataset. Each application has specific features that are mentioned in issues or reviews. So, bug reports in the similar apps can have more similarity to each other than feature requests. Therefore, the better performance of the Rev+(Sim)I compared to the Rev+(S)I for bug reports and the opposite results in the feature requests can be rationalized.
@@
5.2. RQ2: Influences of Dataset Augmentation on Increasing the Generalizability of Model
In RQ2, we investigate the impact of augmenting primary datasets on increasing the generalizability of the trained classifiers. Devine et al. (Devine et al., 2023) demonstrated that models trained on seven labeled datasets do not perform well in predicting unseen ones. Therefore, this experiment demonstrates that augmenting review datasets using auxiliary datasets extracted from labeled issues can contribute to the enhancement of performance in trained classifiers. To this end, we use five primary datasets (Table 10) for training baseline and augmented models. Devine et al. (Devine et al., 2023) introduced two types of models, namely “Single-Dataset” and “Leave-One-Out (LOO)”, for comparison purposes. Their study revealed that the LOO models yield superior results. In this experiment, we also train these two model types using the primary and augmented datasets and compare their results:
  1. (1)
@@ 5.2.1. Results
Table 14. Results of Single-Dataset and LOO Models for Bug Report Classification Using Primary and Augmented Datasets in RQ2 (PD=Primary Dataset, AD=Auxiliary Dataset).
@@
Table 15. Results of Single-Dataset and LOO Models for Feature Request Classification Using Primary and Augmented Datasets in RQ2 (PD=Primary Dataset, AD=Auxiliary Dataset).
@@
Devine et al. (Devine et al., 2023) investigate the performance of review classifiers in predicting new datasets that have not seen any review from them in the training dataset. They demonstrate that models trained on LOO (combined multiple datasets) outperformed Single-Dataset models. In this experiment, LOO models do not necessarily provide the best results but are close to maximum scores. In the bug report, models with the highest F1-scores in order related to augmented PD3 (67.8), augmented PD4 (67.6), baseline PD3 (67.3), and baseline LOO (67.1). In the feature request, models with the highest F1-score in order related to baseline PD4 (50.9), baseline LOO (49.1), and jointly augmented LOO and PD3 (48.8). So, models that augment with the Between-App method can outperform LOO models.
@@
5.3. RQ3: Influences of Volume of Auxiliary Dataset on Model Performance
@@ 5.3.1. Results
Table 16. Results of Bug Report Classification Using Augmented Datasets with Varied Volumes in RQ3 (AD=Auxiliary Dataset).
@@
Table 17. Results of Feature Request Classification Using Augmented Datasets with Varied Volumes in RQ3 (AD=Auxiliary Dataset).
@@
Figure 9. The Average F1-score and Recall for the Bug Report/Feature Request Classifiers Trained on Datasets Augmented with Various Volume Ratios.
@@
5.4. RQ4: Influences of Dataset Augmentation on Transformer-Based Models
In this research question, our goal is to explore the impact of dataset augmentation on the performance of classifiers trained using various transformer-based algorithms. We discuss the results of review classifiers based on four models BERT (ber, 2023), RoBERTa (rob, 2023), ALBERT (alb, 2023), and DistilBERT (dis, 2023). RoBERTa (Liu et al., 2019) stands as “Robustly-Optimized BERT” that performs better than BERT (Devlin et al., 2019) by applying modifications in training and masking strategies. Also, ALBERT (Lan et al., 2019) and DistilBERT (Sanh et al., 2019) use techniques to provide faster and lighter models without significant performance degradation (Izadi et al., 2021). Since transformer-based models outperformed traditional algorithms for our task in previous studies (Hadi and Fard, 2021; Henao et al., 2021), we do not consider them in this experiment. The results of this experiment will facilitate the selection of the appropriate algorithm for training binary review classifiers.
5.4.1. Results
@@
Table 18. Results of Review Classification Using Primary (LOO) and Augmented Datasets with Different Transformer-Based Models in RQ4 (*minimum time)
@@
5.5. Discussion
In this paper, our goal is to extract valuable information from the labeled issues and use them to enrich the training dataset. Our experiments show that using the processed issues as an auxiliary dataset result in an improved review classifier, notably in terms of F1-score and recall metrics. User reviews, especially those related to the bug report, can help developers in maintenance decisions (Gu and Kim, 2015). Therefore, losing a smaller percentage of user reviews related to app maintenance enhances the reliability of the review classifier. Also, since reviews may change over time (Dąbrowski et al., 2022), using processed issues as part of the training dataset makes it easier and faster to update it.
@@ 5.5. Discussion
5.5.1. Training App-Specific Classifiers
@@
5.5.2. Training General-Purpose Classifiers
@@ 6. Related Work
6.1. Approaches Based on App Store Reviews
@@
Some previous research analyzes user reviews based on Natural Language Processing (NLP) and Information Retrieval (IR) techniques. As one of the first research in user review analysis, Iacob et al. (Iacob and Harrison, 2013) propose the MARA tool. They conduct manual analysis and define 237 linguistic rules for extracting feature requests. They rank identified requests by their length and frequency. Guzman et al. (Guzman and Maalej, 2014) automatically analyze user reviews and identify app features at two levels. They filter out the extracted collocations to discover low-level features and compute sentiment scores for them. These features are then organized into groups using topic modeling techniques. Gao et al. (Gao et al., 2015) propose an approach that automatically extracts and prioritizes the main topics derived from user reviews. In addition to the information provided by users, they also use the relevance of the topics to rank them. Also, they visualize and analyze topic dynamics over time. While these unsupervised approaches aid in extracting valuable insights from user reviews, it’s important to note that supervised algorithms can provide better performance in identifying reviews related to predefined classes.
@@
Many studies used traditional learning algorithms to train user review classifiers. Chen et al. (Chen et al., 2014) present AR-Miner, an approach to identify informative user reviews using the EMNB algorithm and group them by applying the topic modeling. They provide a scheme for ranking generated groups and their instances. Villarroel et al. (Villarroel et al., 2016) (Scalabrino et al., 2019) present the CLAP tool that helps developers plan the next release. They classify user reviews in the bug report, suggestion for new feature, and other classes using their constituent words, N-grams, and ratings. Also, they cluster the reviews related to the two first classes, which are of developer interest. Then, clusters were prioritized to recommend the most important ones by defining several independent features. Maalej et al. (Maalej and Nabil, 2015)(Maalej et al., 2016) classify reviews into four classes bug report, feature request, user experience, and rating. They employ this approach based on constituent words, collocations, general metadata, and sentiment scores derived from user reviews. Panichella et al. (Panichella et al., 2015) classify review sentences using a combination of three features: constituent terms, grammatical patterns, and sentiment scores. They analyze developer discussions about maintenance tasks and adapt them with user review categories provided by Pagano et al. (Pagano and Maalej, 2013). As a result, they select helpful categories from the developer’s perspective relevant to the app evolution. Guzman et al. (Guzman et al., 2015) identify a taxonomy of seven categories (bug report, feature strength, feature shortcoming, user request, praise, complaint, and usage scenario) that provide helpful information for the app evolution. They represent the vectorized form of a user review by the tf-idf weighted vector and several numerical features (such as the word and character count). Also, they use four traditional algorithms and their ensembles to classify user reviews based on their taxonomy. Jha et al. (Jha and Mahmoud, 2018) present an approach to classify and summarize helpful reviews. In this approach, reviews represent by semantic frames that are automatically extracted and more efficient than linguistic patterns obtained by manual analysis. Gu et al. (Gu and Kim, 2015) aim to summarize user reviews. First, they classify reviews into five categories, aspect evaluation, bug report, feature request, praise, and other, and use two types of textual features, lexical and structural. Also, they summarize the sentences of the aspect evaluation category by identifying and clustering the aspects and their related opinions and sentiments. This approach employs defined patterns to extract app aspects.
@@
Recent studies show that deep learning algorithms can outperform traditional algorithms in review classification. Aslam et al. (Aslam et al., 2020) use a CNN-based network that considers the textual, non-textual and sentiment features extracted from app reviews as classifier inputs. Also, Stanik et al. (Stanik et al., 2019) present an approach to classify user reviews and tweets in two languages, English and Italian. In addition to traditional models, they use a CNN-based network for classifier training. Due to the relatively small volume of labeled datasets, transfer learning can be helpful in review classification task (Henao et al., 2021). Henao et al. (Henao et al., 2021) argue that BERT-based models outperform models based on traditional machine learning algorithms. Also, Araujo et al. (Araujo et al., 2022) compare the various techniques available for review representation and demonstrate that the pre-trained transformer-based models achieve better results in review analysis tasks. Hadi et al. (Hadi and Fard, 2021) conduct a more comprehensive study comparing pre-trained models and previous approaches for the review classification task in different settings. They focused their research on six labeled datasets from prior works and integrated them and four popular approaches. In addition, they pre-train new models specific to this domain using collected user reviews and show that using them results in better performance. So, we also use pre-trained models in our approach. In most previous works, predictive models have been trained and tested on one dataset. Devine et al. (Devine et al., 2023) discuss how a review classifier trained on one dataset performs well in predicting the intention of reviews related to other apps or datasets. This study uses seven datasets labeled in previous studies and shows that the results of model evaluations on unseen datasets cannot perform as well as the model evaluations on the same dataset. Also, they argue that merging labeled datasets can help overcome this generalizability problem. Our paper provides another solution by augmenting the labeled review dataset with processed issues that can result in better performance than the merged dataset.
6.2. Approaches Based on Other Sources of data
@@
Different sources, such as issue tracker systems, developer emails, and social platforms, contain valuable information for development and maintenance tasks (Devine et al., 2023). Twitter is one of the sources that has been the focus of researchers. Guzman et al. (Guzman et al., 2017) manually analyze tweets related to software applications (including desktop, mobile, or both) and identify 26 categories for their content. Also, they train traditional classifiers to predict the relevance of these categories with the various types of stakeholders. Nayebi et al. (Nayebi et al., 2018) classify tweets and user reviews into three categories: bug reports, feature requests, and others. Then, they extract the topics of each category and analyze their similarity in reviews and tweets. They found that analyzing tweets can provide developers with worthy information besides information extracted from reviews. Sorbo et al. (Di Sorbo et al., 2015) classify the information provided in the developer emails into six categories with manual analysis. They define linguistic patterns to identify the corresponding category. Shi et al. (Shi et al., 2021) present an approach to identify feature requests in developer emails. They categorize sentences using 81 fuzzy rules and find feature requests using sequential pattern mining. In these approaches, training and testing datasets are selected from the same source.
@@
Some studies use available data from one source to help analyze and extract information from another source. Huang et al. (Huang et al., 2018) examine whether the taxonomy introduced by Sorbo et al. (Di Sorbo et al., 2015), provided by analyzing the email contents, can be useful for categorizing GitHub Comments. They customize it and make it suitable for the analysis of GitHub comments. Also, they use a CNN-based neural network to categorize them automatically. Their results show that using classifiers trained on a dataset related to one source (developer email/issues) to predict a dataset related to another source (issues/developer email) does not provide better results than the classifiers that use data from the same source for training and prediction. Therefore, since data from different sources usually have a different nature, we consider the size of the auxiliary datasets to be smaller than the primary datasets in our approach to avoid the excessive biasing model to them. Stanik (Stanik et al., 2019), Henao (Henao et al., 2021), and Devine (Devine et al., 2023) also use Twitter datasets in their studies. However, only the last one examines merging them with review datasets to improve the performance of trained models.
@@
Al-Safoury et al. (Al-Safoury et al., 2022) aggregate the information provided in user reviews and GitHub issues. They extract bug reports and feature requests from them and merge similar ones to create a complete change feature list by computing textual similarity. In our proposed approach, we process issues to adapt with app store reviews. Then, we use them to augment the labeled review dataset. As far as we know, GitHub issues have not been used in previous approaches to improve the performance of the review classification. Zhang et al. (Zhang et al., 2019) utilized the similarity between user reviews and GitHub issues to categorize unlabeled issues. They applied the SURF tool (Sorbo et al., 2016) to identify app reviews related to problem discovery and feature requests and incorporated their similarity with unlabeled issues in the final similarity measure. In this study, our objective differs from their research, as we utilize labeled issues to assist in categorizing user reviews.
@@
Additionally, some studies augment the text of reviews to improve the accuracy of review prediction. Wang et al. (Wang et al., 2019) emphasize that using other sources can help prepare accurate classifiers. They extract feature words related to each category from app changelogs and use them to augment reviews. Liu et al. (Liu, 2018) also utilize app descriptions for this purpose. Our proposed approach differs from this research because we aim to augment the labeled dataset and not add additional information to the text of reviews.
7. Threats to Validity